OpenStudy (anonymous):

solve using the substitution method. 11x + 10y = 77 9x + y = 63

OpenStudy (anonymous):

\[\begin{cases}11x+10y=77\\9x+y=63\end{cases}\] \[\begin{cases}11x+10(63-9x)=77\\y=63-9x\end{cases}\] now find x in first equation and plug it into second

OpenStudy (lilg132):

make one of the variables the same lets make y the same value so 9x + y = 63 (multiply by 10) 90x + 10y = 630 11x + 10y = 77 79x = 553 x = 7 now we know the value of x we use it to work out the value of y

OpenStudy (lilg132):

so 11(7) + 10y = 77 77 + 10y = 77 10y = 77 - 77 10y = 0 10y / 10 = 0 / 10 y = 0 lol

OpenStudy (anonymous):

@lilg132 I think it's not substitution method or it is?

OpenStudy (lilg132):

elimination sorry

OpenStudy (lilg132):

let me do it again using substitution

OpenStudy (anonymous):

I have done most of it, now you need just to find x and y and he should be able to do it himself

OpenStudy (lilg132):

11x + 10y = 77 9x + y = 63 y = 63 - 9x so we use this in the first equation 11x + 10(63 - 9x) = 77 11x + 630 - 90x = 77 11x - 90x = 77 - 630 -79x = -553 -x = -7 x = 7 use this to work out y in the second equation

OpenStudy (lilg132):

9(7) + y = 63 63 + y = 63 y = 63 - 63 y = 0